in a nuclear reactor. Its half-life is approxillaiui that should be used in (1.1) for this isotope? 1.3 In 1947 a large collecti on of papyrus scrolls, including the oldest known manuscript version of portions of the Old Testament, was found in a cave near the Dead Sea; they have come to be known as the Dead Sea Scrolls. The scroll containing the book of e it was found to contain between 75% and 77% of the initial level of carbon 14 . Between which dates was the scroll written? l tobla hanas on the wall of the castle in Winchester. Many would like
Answer #1

me毛) Contul ns only a fracfim where o.15 0.77 our cu m is知finds. ヲ b--K(1994-8) 2 he In p 1216x10千 when t = 6-75 s= 1994-2365